# Projective reflection groups of finite covolume

### v1

@@ 5.2. Finding a bigger face
Let $P$ be a Coxeter polytope with facet set $S$, and let $S^{\prime}\subset S$. We say that $S^{\prime}$ defines a face of $P$ if one of the equivalent conditions below is satisfied:
  1. (i)
@@
[…]
  3. (iii)
@@
Given a subset $T$ of $S$ and a Cartan matrix $A=(A_{st})_{s,t\in S}$ on $S$, we let $A_{T}\coloneqq(A_{st})_{s,t\in T}$. Notice that $A_{T}$ is also a Cartan matrix for any $T\subset S$. Like for the faces of $P$, we qualify $T$ by an adjective if this adjective applies to $A_{T}$. We let $T^{+}$ (resp. $T^{0}$, resp. $T^{-}$) be the union of the irreducible components of $T$ of positive (resp. zero, resp. negative) type. The following lemma is inspired by [Vin71, Thm. 4].
@@
Let $P$ be a Coxeter polytope with facet set $S$ and $T_{1},T_{2}\subset S$ disjoint. If $T_{1}\cup T_{2}$ defines a face of $P$ and $T_{1}\subset T_{2}^{\perp}$, then each of $T_{1}\cup T_{2}^{0}$, $T_{1}\cup T_{2}^{0}\cup T_{2}^{+}$, and $T_{1}\cup T_{2}^{0}\cup T_{2}^{-}$ also defines a face of $P$.
@@
Let $A\coloneqq A_{P}$, and for any $S^{\prime}\subset S$, let $X_{S^{\prime}}\coloneqq(X_{s})_{s\in S^{\prime}}$. Note that we have $X_{T_{1}\cup T_{2}}A_{T_{1}\cup T_{2}}=0$ from the previous paragraph. This decomposes into $X_{T_{1}}A_{T_{1}}=0$ and $X_{T_{2}}A_{T_{2}}=0$, as $T_{1}$ and $T_{2}$ are orthogonal, and similarly, the second identity yields $X_{T}A_{T}=0$ and $X_{T^{\prime}}A_{T^{\prime}}=0$. In particular, we have $X_{I}A_{I}=0$ for every irreducible component $I$ of $T^{\prime}$, but since $I$ is either of positive or negative type, Theorem 3.11 applied to the transpose of $A_{I}$ yields $X_{I}=0$. Hence $X_{T^{\prime}}=0$. ∎
@@ 5.3. Maximal boundary faces
A boundary face of a Coxeter polytope of negative type $P$ is a face of $P$ that lies in $\partial\Omega_{P}$. We order boundary faces using the inclusion of the closures: $f$ is smaller than $g$ if $\overline{f}\subset\overline{g}$. A maximal boundary face of $P$ is a boundary face of $P$ that is maximal. Recall from Theorem 3.8 that a face $f$ of $P$ is a boundary face if and only if the standard subgroup $\Gamma_{f}$ is infinite. Hence $P$ is quasiperfect if and only if every maximal boundary face of $P$ is a parabolic vertex.
@@
Let $P$ be a Coxeter polytope of negative type. A boundary face $f$ of $P$ is maximal if and only if $f$ is perfect. In particular, the maximal boundary faces of $P$ are exactly its parabolic and irreducible loxodromic faces.
@@
Suppose $f$ is maximal and $P_{f}$ has a vertex $v$ in the boundary of its Vinberg domain $\Omega_{f}$. Then the preimage of $v$ under the map $\operatorname{\mathbb{P}}(V)\smallsetminus\operatorname{\mathbb{P}}(V/\operatorname{Span}f)\to\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ is a boundary face of $P$ whose closure contains $f$, which is a contradiction. To prove the converse, suppose $f$ is perfect. Hence $P_{f}\subset\Omega_{f}$ by Theorem 3.8, so $f$ is not properly contained in the closure of any boundary face of $P$. ∎
@@
Let $P$ be a Coxeter polytope of negative type that does not have proper negative type faces. Then the maximal boundary faces of $P$ are its parabolic faces.
@@ 6.1. Facets of boundary faces
Let $P$ be a Coxeter polytope of negative type and $f$ a boundary face of $P$ of dimension at least $1$. Then $\overline{f}$ is a projective polytope in $\operatorname{\mathbb{P}}(\operatorname{Span}f)$, and we let $\Sigma(f)$ denote its facet set. Each $g\in\Sigma(f)$ is also a face of $P$, so we can consider the set $S_{g}\subset S$ of facets of $P$ containing $g$.
@@
Choose an element $s_{g}\in S_{g}\smallsetminus S_{f}$ for each $g\in\Sigma(f)$. Then $\overline{f}$ is the intersection of $\operatorname{\mathbb{P}}(\operatorname{Span}f)$ with
$\bigcap_{g\in\Sigma(f)}\operatorname{\mathbb{P}}\{v\in V\mid\alpha_{s_{g}}(v)\leq 0\}$
In particular, the intersection of $\operatorname{Span}f$ and $\bigcap_{g\in\Sigma(f)}\ker\alpha_{s_{g}}$ must be trivial.
@@
and $S_{2}\coloneqq S_{f}\cup(S_{f}^{\perp}\cap S_{g})\smallsetminus S_{1}$. Then $S_{g}=S_{1}\cup(S_{2}\cup T)$. To get the desired conclusion, we show that (i) $S_{2}\cup T$is of negative type, and (ii) $S_{1}$is empty if $f$ is a maximal boundary face of $P$.
@@
    Let $I$ be an irreducible component of $S_{2}\cup T$. If $I\subset S_{2}$, then since $S_{2}$ is a union of irreducible components of $S_{f}\cup(S_{f}^{\perp}\cap S_{g})$, $I$ is an irreducible component of $S_{f}\cup(S_{f}^{\perp}\cap S_{g})$. Moreover, $I$ is also orthogonal to $(S_{2}\cup T)\smallsetminus I\supset T$, contradicting the definition of $S_{2}$. On the other hand, if $I\subset T$, then $I$ is orthogonal to $(S_{2}\cup T)\smallsetminus I$, which contains $S_{2}$. Since $S_{1}$ and $S_{2}\cup T$ are orthogonal, $I$ is also orthogonal to $S_{1}\cup S_{2}\supset S_{f}$, so $I\subset S_{f}^{\perp}\cap S_{g}\subset S_{2}$, which is a contradiction. Hence $I\cap S_{2}\neq\varnothing$ and $I\cap T\neq\varnothing$.
    We now claim that $I\cap S_{2}$ is a union of irreducible components of $S_{2}$. Let $J$ be an irreducible component of $S_{2}$ that intersects $I$. Since $I$ is an irreducible component of $S_{2}\cup T\supset J$, $J\cap I$ and $J\smallsetminus I$ are orthogonal, so we must have $J\smallsetminus I=\varnothing$, i.e. $J\subset I$.
    In particular, $I\cap S_{2}$ has no irreducible components of positive type as $S_{f}$ is either of zero or negative type and $S_{f}^{\perp}\cap S_{g}$ is of negative type. Hence $W_{I\cap S_{2}}$ is a proper standard subgroup of $W_{I}$ that is not spherical, and $W_{I}$ is large by Theorem 3.5 and Remark 3.6. So II is of negative type by Proposition 4.5.
@@
    Suppose $f$ is a maximal boundary face of $P$, and $S_{1}\neq\varnothing$. Then $S_{g}=S_{1}\cup(S_{2}\cup T)$ is a nontrivial orthogonal decomposition and $S_{2}\cup T$ is of negative type by (i), so there exists a face $f^{\prime}$ of $P$ such that $S_{f^{\prime}}=S_{1}$ by Lemma 5.3. Since $S_{1}\neq\varnothing$ is not of positive type and $S_{2}\cap T\neq\varnothing$, $f^{\prime}$ is a boundary face of $P$ that properly contains $f$, which is a contradiction. Hence $S_{1}=\varnothing$.
Therefore, if $f$ is a maximal boundary face, then $S_{g}=S_{2}\cup T$ is of negative type by (i) and (ii). On the other hand, if $f$ is of negative type, then $S_{1}$ is also of negative type as a union of irreducible components of $S_{f}\cup(S_{f}^{\perp}\cap S_{g})$, so $S_{g}=S_{1}\cup(S_{2}\cup T)$ is of negative type by (i) as well. ∎
@@
Suppose $P$ has a parabolic face $f$ of positive dimension. Since $P$ has no proper negative type face, Proposition 6.3 implies that $S_{g}\cap S_{f}^{\perp}\neq\varnothing$ for all $g\in\Sigma(f)$. Hence we can choose a facet $s_{g}\in S_{g}\cap S_{f}^{\perp}$ for each $g\in\Sigma(f)$.
By Remark 3.12, we can find a vector $\Lambda=(\lambda_{s})_{s\in S_{f}}>0$ such that $A_{f}\Lambda=0$. If we let $x\coloneqq\sum_{s\in S_{f}}\lambda_{s}v_{s}$, we see that $\alpha_{s}(x)=0$ for all $s\in S_{f}$, i.e. $x\in\operatorname{Span}f$. Notice also that if we let $\lambda_{s}=0$ for $s\in S\smallsetminus S_{f}$ and $\Lambda^{\prime}=(\lambda_{s})_{s\in S}\geq 0$, then $x=0$ implies that $A_{P}\Lambda^{\prime}=0\geq 0$, so by Theorem 3.11, we must have $\Lambda^{\prime}=0$, which is a contradiction.
Thus $x$ is a nonzero element of $\operatorname{Span}f\cap\operatorname{Span}(v_{s})_{s\in S_{f}}$. However, notice that $\operatorname{Span}(v_{s})_{s\in S_{f}}\subset\bigcap_{g\in\Sigma(f)}\ker\alpha_{s_{g}}$ since $\{s_{g}\mid g\in\Sigma_{f}\}\subset S_{f}^{\perp}$. It follows that the intersection of $\operatorname{Span}f$ with $\bigcap_{g\in\Sigma(f)}\ker\alpha_{s_{g}}$ is nonzero, contradicting Remark 6.2.
Therefore, $P$ has no parabolic face of positive dimension. Since all non-positive type faces of $P$ are boundary faces, and are therefore contained in a parabolic face of $P$ by Proposition 5.4, we conclude that the proper faces of $P$ are either parabolic vertices, or elliptic. Hence $P$ is quasiperfect. ∎
@@
By Corollary 5.2, a finite volume Coxeter polytope of negative type $P$ has no proper face of negative type. By Proposition 6.4, a Coxeter polytope of negative type that has no proper negative type faces is quasiperfect. The last implication is given by [Mar17, Thm. 6.3], since a quasiperfect Coxeter polytope of negative type must be loxodromic irreducible by Proposition 4.9. ∎
@@ 7. Minimality of the Vinberg domain
We now inquire about when a reflection group of $\operatorname{SL^{\pm}}(V)$ preserves a unique properly convex domain in $\operatorname{\mathbb{P}}(V)$. This along with the following results generalize [Mar17, Thm. 8.2].
@@
Let $P\subset\operatorname{\mathbb{P}}(V)$ be a Coxeter polytope of negative type with facets $S$. Then $\Omega_{P}$ is the unique $\Gamma_{P}$-invariant properly convex domain in $\operatorname{\mathbb{P}}(V)$ if and only if $P$ is quasiperfect and $|S|\geq 3$.
In order to prove Theorem 7.1, we start by finding when the candidates for the largest and the smallest properly convex domains are equal, i.e. when the Vinberg domain equals the interior of the convex hull of the proximal limit set.
Recall that by convention, the convex hull of a subset $X$ of the closure of a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(V)$ is taken to mean the convex hull of $X$ within any affine chart of $\operatorname{\mathbb{P}}(V)$ containing $\Omega$.
Theorem 7.2.
Let $P$ be an irreducible Coxeter polytope of negative type. Then $P$ is quasiperfect if and only if $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$.
@@
Suppose first that $P$ is an irreducible quasiperfect Coxeter polytope of negative type. $P$ is irreducible loxodromic by Proposition 4.9, and is thus loxodromic $2$-perfect of finite volume in $\Omega_{P}$ by Theorem 6.1. It follows that $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$ by [Mar17, Thm. 8.2] and its proof.
For the converse, we start by assuming that $P$ is an irreducible Coxeter polytope of negative type such that $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$. By Proposition 3.14, $\Lambda_{P}\subset\operatorname{\mathbb{P}}(\operatorname{Span}(v_{s})_{s\in S})$ is non-empty, and since $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$ by assumption, we have $\operatorname{Span}(v_{s})_{s\in S}=V$. In the case where $\Gamma_{P}$ is affine of type $\tilde{A}_{d}$, $P$ is perfect, so it has finite volume in its Vinberg domain. So by Proposition 4.5, we can suppose that $\Gamma_{P}$ is large. In that case, Proposition 4.11 and Remark 4.12 imply that $P\subset P\cap\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$, so that $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$.
Under this assumption, we can show the following claim, whose rather long proof we delay for the sake of the clarity of the proof of this theorem.
Claim 7.2.1.
If $f\subsetneq P$ is a face of negative type that is the join of minimal negative type faces of $P$, then there is a larger face $g\supsetneq f$ of $P$ that is the join of $f$ with a minimal negative type face of $P$.
If we suppose that $P$ has proper faces of negative type, then there are minimal proper faces of negative type, and we can consider $f\subset P$ a face of $P$ that is maximal for inclusion among the faces of $P$ that are the join of minimal negative type faces of $P$. By Claim 7.2.1, $f$ must be equal to $P$ itself by maximality, hence $P$ is the join of its minimal negative type faces. Thus $P$ is decomposable, contradicting irreducibility. Therefore $P$ has no proper faces of negative type, so it is quasiperfect by Theorem 6.1. ∎
Proof of Claim 7.2.1.
We start by making the following observation about proper (minimal) negative type faces of $P$.
Claim 7.2.2.
@@
If $f$ is a proper negative type face of $P$, then $f\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S_{f}^{\perp}})$. If moreover $f$ is a minimal negative type face of $P$, then $f$ can be endowed with a quasiperfect Coxeter polytope structure of negative type $(\alpha_{s},v_{s})_{s\in S_{f}^{\perp}}$.
Let $f$ be the join of minimal negative type faces $(f_{i})_{i=1}^{k}$ of $P$. We consider
$Q\coloneqq\bigcap_{s\in S_{f}\cup S_{f}^{\perp}}\operatorname{\mathbb{P}}\{\alpha_{s}\leq 0\}$
Observe that $Q$ is nonempty and properly convex as $\bigcap_{s\in S_{f}\cup S_{f}^{\perp}}\{\alpha_{s}=0\}=\varnothing$ by Remark 6.2. Indeed, the facets of $f$ are determined by the elements of $S_{f}^{\perp}$ since $f$ is the join of the $f_{i}$ and for each $i$, the facets of $f_{i}$ are given by $S_{f_{i}}^{\perp}\subset S_{f}^{\perp}$ by Claim 7.2.2. Moreover, we can endow $Q$ with a Coxeter polytope structure with the data $(\alpha_{s},v_{s})_{s\in S_{f}\cup S_{f}^{\perp}}$.
By Lemma 5.3 applied to the polytope $Q$ with $T_{1}=S_{f}^{\perp}$ and $T_{2}=S_{f}$, $S_{f}^{\perp}$ defines a face $f_{Q}^{\perp}$ of $Q$. Since the facets of $f$ are determined by the elements of $S_{f_{1}}^{\perp}\cup\dots\cup S_{f_{k}}^{\perp}\subset S_{f}^{\perp}$ which is a subset of the set of facets of $Q$, and $P\subset Q$, $f$ is also a face of $Q$. It follows that $Q$ is the join of $f$ and $f_{Q}^{\perp}$.
Since $f$ is a negative type face of $P$, $S_{f}$ is of negative type, so we can find a vector $\Lambda\coloneqq(\lambda_{s})_{s\in S_{f}}>0$ such that $A_{S_{f}}\Lambda<0$ by Remark 3.12. In particular, if we let $x=\sum_{s\in S_{f}}\lambda_{s}v_{s}$, we have that $\alpha_{t}(x)\leq 0$ for all $t\in S$ with equality if and only if $t\in S_{f}^{\perp}$. Hence $S_{f}^{\perp}$ also defines a face $f^{\perp}\coloneqq P\cap f_{Q}^{\perp}$ of $P$ such that $S_{f^{\perp}}=S_{f}^{\perp}$.
We claim that $S_{f}^{\perp}=S_{f_{1}}^{\perp}\sqcup\dots\sqcup S_{f_{k}}^{\perp}$. We start by observing that the $S_{f_{i}}^{\perp}$ are pairwise disjoint as $S_{f_{i}}^{\perp}\subset S_{f_{j}}$ whenever $j\neq i$ (since $f$ is the join of the $f_{i}$ whose facets are given by $S_{f_{i}}^{\perp}$), so the right-hand side is indeed a disjoint union, and is clearly included in $S_{f}^{\perp}$. Moreover, observe that the hyperplane spanned by a facet in $S_{f}^{\perp}=S_{f^{\perp}}$ must define a facet of $f$ since it already contains $\operatorname{\mathbb{P}}(\operatorname{Span}f^{\perp})$, and since the facets of $f$ are exactly given by the $S_{f_{i}}^{\perp}$, we have the reverse inclusion.
Now, since $f$ is the join of finitely many Coxeter polytopes of negative type whose Coxeter structure is the restriction of the Cartan structure of $P$, $f$ is also a Coxeter polytope of negative type with the Coxeter structure given by the restriction of that of $P$ to $S_{f}^{\perp}$. Therefore, $A_{S_{f}^{\perp}}$ is of negative type, so $f^{\perp}$ is a negative type face of $P$, and we can find a minimal negative type face $f^{\prime}$ of $P$ contained in $f^{\perp}$ (we may have $f^{\prime}=f^{\perp}$ if $f^{\perp}$ itself is minimal).
Observe that $S_{f}^{\perp\perp}=S_{f}$ in $S$. This is because $s\in S_{f}^{\perp\perp}$ if and only if $\alpha_{s}(v_{t})=0$ for all $t\in S_{f}^{\perp}$, if and only if $f\subset\operatorname{\mathbb{P}}(\ker\alpha_{s})$ as $\operatorname{Span}f=\operatorname{Span}(v_{s})_{s\in S_{f}^{\perp}}$ by Claim 7.2.2, if and only if $s\in S_{f}$. In particular, we have $S_{f^{\perp}}=S_{f}^{\perp}\subset S_{f^{\prime}}$, so that $S_{f^{\prime}}^{\perp}\subset S_{f}^{\perp\perp}=S_{f}$. By Claim 7.2.2, we see that the facets of $f^{\prime}$ are determined by elements of $S_{f^{\prime}}^{\perp}\subset S_{f}$, so $f^{\prime}$ is also a face of $Q$. Since $Q$ is the join of $f_{Q}^{\perp}$ and $f$, and $f^{\prime}$ is a face of $f_{Q}^{\perp}$, the relative interior $g$ of the convex hull of $f^{\prime}$ and $f$ is a face of $Q$, which must be included in $P$ since both $f$ and $f^{\prime}$ are. However, we also have that $P\subset Q$, so $P\cap\operatorname{\mathbb{P}}(\operatorname{Span}g)=Q\cap\operatorname{\mathbb{P}}(\operatorname{Span}g)=g$ is a face of $P$ containing $f$ that is the join of minimal negative type faces of $P$. ∎
Proof of Claim 7.2.2.

### v2

@@ 5.2. Finding a bigger face
Recall the following definition from [Vin71].
Definition 5.3.
Let $P\subset\operatorname{\mathbb{P}}(V)$ be a Coxeter polytope with facet set $S$. We say that a subset $S^{\prime}$ of $S$ defines a face of $P$ if one of the equivalent conditions below is satisfied:
  1. (1)
@@
    $\begin{cases}\alpha_{s}=0&\text{for $s\in S^{\prime}$}\\ \alpha_{s}<0&\text{for $s\not\in S^{\prime}$}\end{cases}$
  3. (3)
@@
Given a Cartan matrix $A=(A_{st})_{s,t\in S}$ on $S$ and a subset $T$ of $S$, recall that $A_{T}\coloneqq(A_{st})_{s,t\in T}$ is also a Cartan matrix. Like for the faces of $P$, we qualify $T$ by an adjective if this adjective applies to $A_{T}$. We let $T^{+}$ (resp. $T^{0}$, resp. $T^{-}$) be the union of the irreducible components of $T$ that are of positive (resp. zero, resp. negative) type. The following lemma is inspired by [Vin71, Thm. 4].
@@
Let $P$ be a Coxeter polytope with facet set $S$ and $T_{1},T_{2}\subset S$ disjoint. If $T_{1}\cup T_{2}$ defines a face of $P$ and $T_{1}\subset T_{2}^{\perp}$, then $T_{1}\cup T_{2}^{0}$, $T_{1}\cup T_{2}^{0}\cup T_{2}^{+}$, and $T_{1}\cup T_{2}^{0}\cup T_{2}^{-}$ also define a face of $P$.
@@
Let $A\coloneqq A_{P}$, and for any $S^{\prime}\subset S$, let $X_{S^{\prime}}\coloneqq(X_{s})_{s\in S^{\prime}}$. Note that we have $X_{T_{1}\cup T_{2}}A_{T_{1}\cup T_{2}}=0$ from the previous paragraph. This decomposes into $X_{T_{1}}A_{T_{1}}=0$ and $X_{T_{2}}A_{T_{2}}=0$, as $T_{1}$ and $T_{2}$ are orthogonal, and similarly, the second identity yields $X_{T}A_{T}=0$ and $X_{T^{\prime}}A_{T^{\prime}}=0$. In particular, we have $X_{I}A_{I}=0$ for every irreducible component $I$ of $T^{\prime}$, but since $A_{I}$ is either of positive or negative type, Fact 3.11 applied to the transpose of $A_{I}$ yields $X_{I}=0$. Hence $X_{T^{\prime}}=0$. ∎
@@ 5.3. Maximal boundary faces
A boundary face of a Coxeter polytope of negative type $P$ is a face of $P$ that lies in $\partial\Omega_{P}$. A maximal boundary face of $P$ is a boundary face of $P$ that is maximal. Recall from Fact 3.8 that a face $f$ of $P$ is a boundary face if and only if the standard subgroup $\Gamma_{f}$ is infinite.
@@
Let $P$ be a Coxeter polytope of negative type. A boundary face $f$ of $P$ is maximal if and only if $f$ is perfect. In particular, the maximal boundary faces of $P$ are exactly its parabolic and perfect loxodromic faces.
@@
Suppose $f$ is maximal and $P_{f}$ has a vertex $v$ in the boundary of its Vinberg domain $\Omega_{f}$. Then the preimage of $v$ under the map $\operatorname{\mathbb{P}}(V)\smallsetminus\operatorname{\mathbb{P}}(\operatorname{Span}f)\to\operatorname{\mathbb{P}}(V/\operatorname{Span}f)$ is a boundary face of $P$ whose closure contains $f$, which is a contradiction. To prove the converse, suppose $f$ is perfect. Then $P_{f}\subset\Omega_{f}$ by Fact 3.8, so $f$ is not properly contained in the closure of any boundary face of $P$. ∎
@@
Let $P$ be a Coxeter polytope of negative type that has no proper negative type faces. Then the maximal boundary faces of $P$ are its parabolic faces.
@@ 6.1. Facets of boundary faces
Let $P$ be a Coxeter polytope of negative type and $f$ a boundary face of $P$ of dimension at least $1$. Then $\overline{f}$ is a projective polytope in $\operatorname{\mathbb{P}}(\operatorname{Span}f)$, and we let $\operatorname{\Sigma}(f)$ denote its facet set. Each $g\in\operatorname{\Sigma}(f)$ is also a face of $P$, so we can consider the set $S_{g}\subset S$ of facets of $P$ containing $g$.
@@
Choose an element $s_{g}\in S_{g}\smallsetminus S_{f}$ for each $g\in\operatorname{\Sigma}(f)$. Then $\overline{f}$ is the intersection of $\operatorname{\mathbb{P}}(\operatorname{Span}f)$ with
$\bigcap_{g\in\operatorname{\Sigma}(f)}\operatorname{\mathbb{P}}\{v\in V\mid\alpha_{s_{g}}(v)\leq 0\}$
In particular, the intersection of $\operatorname{Span}f$ and $\bigcap_{g\in\operatorname{\Sigma}(f)}\ker\alpha_{s_{g}}$ must be trivial.
@@
and $S_{2}\coloneqq S_{f}\cup(S_{f}^{\perp}\cap S_{g})\smallsetminus S_{1}$. Note that $S_{g}=S_{1}\cup(S_{2}\cup T)$ and $S_{2}\cup T\subset S_{1}^{\perp}$. To get the desired conclusion, we show that (i) $S_{2}\cup T$is of negative type, and (ii) $S_{1}$is empty if $f$ is a parabolic face of $P$.
@@
    Let $I$ be an irreducible component of $S_{2}\cup T$. If $I\subset S_{2}$, then since $S_{2}$ is a union of irreducible components of $S_{f}\cup(S_{f}^{\perp}\cap S_{g})$, $I$ is an irreducible component of $S_{f}\cup(S_{f}^{\perp}\cap S_{g})$. Moreover, $I$ is also orthogonal to $(S_{2}\cup T)\smallsetminus I\supset T$, so $I\subset S_{1}$ by definition of $S_{1}$, which is a contradiction. On the other hand, if $I\subset T$, then $I$ is orthogonal to $(S_{2}\cup T)\smallsetminus I$, which contains $S_{2}$. Since $S_{1}$ and $S_{2}\cup T$ are orthogonal, it follows that $I$ is orthogonal to $S_{1}\cup S_{2}\supset S_{f}$, so $I\subset S_{f}^{\perp}\cap S_{g}$, which contradicts $I\subset T$. Therefore we have both $I\cap S_{2}\neq\varnothing$ and $I\cap T\neq\varnothing$.
    We now claim that $I\cap S_{2}$ is a union of irreducible components of $S_{2}$. Let $J$ be an irreducible component of $S_{2}$ that intersects $I$. Since $I$ is an irreducible component of $S_{2}\cup T\supset J$, $J\cap I$ and $J\smallsetminus I$ are orthogonal, and since $J$ is irreducible, $J\smallsetminus I=\varnothing$, i.e. $J\subset I$.
    In particular, $I\cap S_{2}$ has no irreducible components of positive type as $S_{f}$ is either of zero or negative type and $S_{f}^{\perp}\cap S_{g}$ is of negative type. Thus $W_{I\cap S_{2}}$ is a proper standard subgroup of $W_{I}$ that is not spherical, so $W_{I}$ is large by Fact 3.5 and Remark 3.6. Hence $I$ is of negative type by Fact 4.5.
@@
    Suppose $f$ is a maximal boundary face of $P$ and $S_{1}\neq\varnothing$. Then $S_{g}=S_{1}\cup(S_{2}\cup T)$ is a nontrivial orthogonal decomposition and $S_{2}\cup T$ is of negative type by (i), so there exists a face $f^{\prime}$ of $P$ such that $S_{f^{\prime}}=S_{1}$ by Lemma 5.4. Since $S_{1}\neq\varnothing$ is not of positive type and $S_{2}\cap T\neq\varnothing$, $f^{\prime}$ is a boundary face of $P$ that properly contains $f$, which is a contradiction. Hence $S_{1}=\varnothing$.
Therefore, if $f$ is a parabolic face, then $S_{g}=S_{2}\cup T$ is of negative type by (i) and (ii). On the other hand, if $f$ is of negative type, then $S_{1}$ is also of negative type as a union of irreducible components of $S_{f}\cup(S_{f}^{\perp}\cap S_{g})$, so $S_{g}=S_{1}\cup(S_{2}\cup T)$ is also of negative type by (i). ∎
@@
Suppose $P$ has a parabolic face $f$ of positive dimension. Since $P$ has no proper negative type faces, Proposition 6.3 implies that $S_{g}\cap S_{f}^{\perp}\neq\varnothing$ for all $g\in\operatorname{\Sigma}(f)$. Hence we can choose a facet $s_{g}\in S_{g}\cap S_{f}^{\perp}$ for each $g\in\operatorname{\Sigma}(f)$.
By Remark 3.12, we can find a vector $\Lambda=(\lambda_{s})_{s\in S_{f}}>0$ such that $A_{f}\Lambda=0$. If we let $x\coloneqq\sum_{s\in S_{f}}\lambda_{s}v_{s}$, we see that $\alpha_{s}(x)=0$ for all $s\in S_{f}$, i.e. $x\in\operatorname{Span}f$. Notice also that if we let $\lambda_{s}=0$ for $s\in S\smallsetminus S_{f}$ and $\Lambda^{\prime}=(\lambda_{s})_{s\in S}\geq 0$, then $x=0$ implies that $A_{P}\Lambda^{\prime}=0\geq 0$, so by Fact 3.11, we must have $\Lambda^{\prime}=0$, which is a contradiction.
Thus $x$ is a nonzero element of $V_{v}\cap\operatorname{Span}f$. However, note that $V_{v}\subset\allowbreak\bigcap_{g\in\operatorname{\Sigma}(f)}\ker\alpha_{s_{g}}$ since $\{s_{g}\mid g\in\operatorname{\Sigma}_{f}\}\subset S_{f}^{\perp}$. Hence the intersection of $\operatorname{Span}f$ with $\bigcap_{g\in\operatorname{\Sigma}(f)}\ker\alpha_{s_{g}}$ is nonzero, contradicting Remark 6.2.
Therefore, $P$ has no parabolic faces of positive dimension. Since all non-positive type faces of $P$ are boundary faces, and are therefore contained in a parabolic face of $P$ by Proposition 5.5, we conclude that the proper faces of $P$ are either parabolic vertices, or elliptic. Hence, $P$ is quasiperfect. ∎
@@
By Corollary 5.2, a finite volume Coxeter polytope of negative type $P$ has no proper faces of negative type. By Proposition 6.4, a Coxeter polytope of negative type that has no proper negative type faces is quasiperfect. The last implication is given by [Mar17, Thm. 6.3], since a quasiperfect Coxeter polytope of negative type must be loxodromic irreducible by Fact 4.10. ∎
@@ 7. Minimality of the Vinberg domain
We now set out to prove Theorem 1.3. Along with Theorem 7.1 below, this provides a generalization of [Mar17, Thm. 8.2]. We start by finding when the candidates for the largest and the smallest properly convex domains are equal, i.e. when the Vinberg domain equals the interior of the convex hull of the proximal limit set.
Recall that by convention, the convex hull of a subset $X$ of the closure of a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(V)$ is taken to mean the convex hull of $X$ within any affine chart of $\operatorname{\mathbb{P}}(V)$ containing $\Omega$. Similarly, “the” join of subsets of the closure of a properly convex domain $\Omega$ is the unique join of these subsets lying in $\overline{\Omega}$.
@@
Let $P$ be a negative type Coxeter polytope. The following are equivalent
  1. (i)
    $\Omega_{P}=\operatorname{int}\operatorname{Conv}(\Lambda_{P})$;
  2. (ii)
    $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$;
  3. (iii)
    $P$ is a join of quasiperfect Coxeter polytopes of negative type.
We will need a few additional lemmas.
Lemma 7.2.
Let $P$ be a perfect Coxeter polytope of negative type. Then
$P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$
@@
Let $p$ be a vertex of $P$. Since $P$ is perfect, it is quasiperfect, so by Fact 4.10, it is indecomposable. In particular, we can find a facet $s_{p}$ of $P$ that is not in $S_{p}\cup S_{p}^{\perp}$. Let $X=(\alpha_{s}(v_{s_{p}}))_{s\in S_{p}}\leq 0$. Notice that $X\neq 0$ as $s_{p}\not\in S_{p}^{\perp}$.
Since $P$ is perfect, $p$ is elliptic, so $A_{S_{p}}$ is of positive type. By Fact 3.11, we can find a vector $\Lambda=(\lambda_{s})_{s\in S_{p}}$ such that $A_{S_{p}}\Lambda=-X$ by splitting $S_{p}$ into its irreducible components. Hence $A_{S_{p}}\Lambda\geq 0$, so $\Lambda\geq 0$ by Fact 3.11 again. If we let $x:=v_{s_{p}}+\sum_{s\in S_{p}}\lambda_{s}v_{s}$, the identity $A_{S_{p}}\Lambda+X=0$ implies that $\alpha_{s}(x)=0$ for all $s\in S_{p}$. Thus $p=[x]\in\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$.
Since this holds for all the vertices of $P$, $P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$. ∎
Lemma 7.3.
Let $P\subset\operatorname{\mathbb{P}}(V)$ be a negative type Coxeter polytope such that
$P\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S})$
If $f$ is a proper negative type face of $P$, then
$f\subset\operatorname{\mathbb{P}}(\operatorname{Conv}(v_{s})_{s\in S_{f}^{\perp}})$
@@
If moreover $f$ is minimal, then $(\alpha_{s},v_{s})_{s\in S_{f}^{\perp}}$ is a quasiperfect Coxeter polytope structure of negative type on $f$ seen as a subset of $\operatorname{\mathbb{P}}(\operatorname{Span}f)$.
Proof.